Please login if you have access to particular applications.
|Subject:||Both ADCPs were found to be off this morning, and had config change to ext trig, they were put back on int trig and restarted at 19:45 30/6/2016 UTC|
|When:|| 2016-06-30 21:29:12 UTC |
- entered into Elog/Everlog as 2016-06-30 21:29:12
Equipment: 75/150 KHz ADCP [view]
|Position:||Latitude: -19.115 Longitude: 175.0006 Position from underway logging|